Session: Oct 2009 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Thu Nov 5 17:10:36 2009) Minor failures : ANT
Due to antenna control problem, We lost 4 scans from 15:10UT to 15:26UT.

Shanghai (Bo Xia at Thu Nov 5 17:37:07 2009) Minor failures :
From UT 12:00 to 13:03, LO problem. After UT 13:03,it is normal.

Onsala (Michael Lindqvist at Thu Nov 5 18:32:10 2009) Minor failures :
Tracking problems between 1305-1337 UTC. Strange Tsys at end.

Onsala (Michael Lindqvist at Fri Nov 6 08:14:59 2009) Minor failures :
Addition: We seems to have problem with the noise diode.

Torun (Kaz Borkowski at Fri Nov 6 08:37:14 2009) Minor failures : RECO
Some data lost due to problems with the recorder which required reset of Mk5 around the third scheduled ftp (15:40 UTC).

Westerbork (Geert Kuper at Sat Nov 7 21:50:51 2009) Success :
Wb used in phased array mode. No known problems.

Effelsberg (Uwe Bach at Mon Nov 9 11:19:56 2009) Success :

Noto (P. Cassaro at Mon Nov 30 09:52:16 2009) Minor failures :
Problems with the MK5 software in extracting the files for the ftp fringe test. Files extracted and sent by hand.

Jodrell_Bank (Alastair Gunn at Thu Dec 3 11:58:39 2009) Success :
The Lovell Telescope was used (Jb1). No known problems.

Cambridge (Alastair Gunn at Thu Dec 3 11:59:07 2009) Success :
The Darnhall telescope was used. No known problems.